OpenStudy (anonymous):

limit of x^3*e^(-x) when x goes to -infinity

OpenStudy (helder_edwin):

\[ \large \lim_{x\to-\infty}x^3e^{-x}=\lim_{x\to-\infty}\frac{x^3}{e^x}=0 \]

OpenStudy (anonymous):

i think the answer is minus infinity, can you explain how you did the last move and got zero ? Thanks !

OpenStudy (helder_edwin):

well \[ \large a^{-n}=\frac{1}{a^n} \]

OpenStudy (helder_edwin):

u r right, sorry, it is \(-\infty\)

OpenStudy (anonymous):

so how you know it ? can you explain ? i can't figure this one out